OpenStudy (anonymous):

can someone help me pleaseeee 5^2+19z-4

OpenStudy (imtant):

Do you want to find z?right?

OpenStudy (anonymous):

yes

OpenStudy (imtant):

If that so,the solution like this..19z=-25+4, then you get 19z=-21,and finally z=-21/19..

OpenStudy (anonymous):

i dont get it o.O

OpenStudy (imtant):

5^2+19z-4=0 (bring 5^2 to the right and change to - value,same goes to -4 that then be +4, like this.. 19z=-25+4 and now solve like it. 19z=-21 and z equal to -21/19..Is that clear?

OpenStudy (anonymous):

so the answer is -21 thats it?

OpenStudy (imtant):

No.it is -21/19..because you need to find z..so 19 must bring to the right..that will become division..ok?

OpenStudy (anonymous):

ok so i would have to divide it or leave it like that ?

OpenStudy (imtant):

you can leave it like that.

OpenStudy (anonymous):

oh ok thank you

Directrix (directrix):

Is the problem supposed to be this: Solve z^2 + 19z - 4 = 0 for z? Just wondering. @naomi_29

OpenStudy (imtant):

Yw..

OpenStudy (anonymous):

i wrote the question wrong it was supposed to be 5z^2+19z-4

Directrix (directrix):

What are the instructions for the problem?

OpenStudy (imtant):

So you need to factorize it..

OpenStudy (imtant):

There will be two answers..that is 1/5 and-4..

OpenStudy (anonymous):

yes i need factors

Directrix (directrix):

>yes i need factors >>that is 1/5 and-4.. These are roots, not factors.

OpenStudy (imtant):

the factors are (5z-1)(z+4)..right?